# DRDT3: Diffusion-Refined Decision Test-Time Training Model

### v1

@@ Abstract
Decision Transformer (DT), a trajectory modeling method, has shown competitive performance compared to traditional offline reinforcement learning (RL) approaches on various classic control tasks. However, it struggles to learn optimal policies from suboptimal, reward-labeled trajectories. In this study, we explore the use of conditional generative modeling to facilitate trajectory stitching given its high-quality data generation ability. Additionally, recent advancements in Recurrent Neural Networks (RNNs) have shown their linear complexity and competitive sequence modeling performance over Transformers. We leverage the Test-Time Training (TTT) layer, an RNN that updates hidden states during testing, to model trajectories in the form of DT. We introduce a unified framework, called Diffusion-Refined Decision TTT (DRDT3), to achieve performance beyond DT models. Specifically, we propose the Decision TTT (DT3) module, which harnesses the sequence modeling strengths of both self-attention and the TTT layer to capture recent contextual information and make coarse action predictions. We further integrate DT3 with the diffusion model using a unified optimization objective. With experiments on multiple tasks of Gym and AntMaze in the D4RL benchmark, our DT3 model without diffusion refinement demonstrates improved performance over standard DT, while DRDT3 further achieves superior results compared to state-of-the-art conventional offline RL and DT-based methods.
Index Terms:
Offline reinforcement learning, Transformer, diffusion model, test-time training.
I Introduction
@@
Reinforcement learning aims to train agents through interaction with the environment and receiving reward feedback. It has been widely applied in areas such as robotics [1, 2], autonomous driving [3, 4], recommendation systems [5], transportation systems [6, 7], and games [8]. Offline RL, also known as batch RL, eliminates the reliance on online interactions with the environment [9] by directly learning policies from offline datasets composed of historical trajectories. Therefore, it is promising for applications where interactions are risky and costly. Directly applying dynamic programming-based online RL methods to offline RL problems typically learns poor-quality policy due to overestimated out-of-distribution actions caused by distribution shift [10]. Conventional offline RL generally addresses such issues by regularizing policy to stay close to behavior policy [11] or constraining values of out-of-distribution actions [12, 13].
Figure 1: Normalized scores of the latest DT variants and our proposed DRDT3 and DT3 on Gym tasks with Medium datasets. DRDT3 and DT3 significantly improve the performance over DT on sub-optimal datasets.
@@
DT [14], an innovative offline RL approach that formulates policy learning as a sequence modeling problem, has shown superior performance over conventional offline RL methods on some benchmarks. By leveraging the strong sequence modeling capacity of Transformer structures [15, 16], DT models trajectories autoregressively, bypassing the necessity for bootstrapping and learning a value function for a given state. Fed with the recent context, a historical subtrajectory composed of states, actions, and return-to-go (RTG) (cumulative rewards from the current step), DT predicts actions autoregressively. However, the abandonment of dynamic programming in DT makes it lack stitching ability [10]. Some recent work studies such problem. For instance, Q-learning DT [10] relabels RTG with value function learned from Q-learning, enhancing DT to learn better policy from sub-optimal datasets. Elastic DT [17] adjusts the context length according to the evaluated optimality of the context to achieve trajectory stitching. Our work instead addresses such a problem using generative modeling.
@@
Recently, diffusion models  [18, 19, 20] have gained recognition for their extraordinary ability to generate high-quality complex data, such as images and texts. Additionally, recent works have also introduced the diffusion model to offline RL [21, 22, 18]. Similar to DT, Diffuser [21] also works as a trajectory generator, using expressive diffusion models instead of Transformers. Unlike existing diffusion-based offline RL methods, our work leverages diffusion models as a refinement tool.
@@
We propose DRDT3, a framework that combines a novel DT-style trajectory modeling method and a conditional diffusion model. Specifically, we first introduce a DT3 module, which harnesses both self-attention and the TTT layer to achieve reduced complexity and improved sequence modeling ability. This module predicts coarse action representations based on recent context, which are then used as prior knowledge or conditions for a denoising diffusion probabilistic model (DDPM) [18]. Additionally, we present a novel gated MLP noise approximator, designed to capture essential information from both noisy input and coarse action predictions, facilitating the denoising process. To integrate the DT3 module and diffusion model effectively, we employ a unified optimization objective with two key components: an action representation term, which ensures that the DT3 module generates coarse action conditions close to the optimal action distribution, and an action refinement term that constrains the diffusion model to sample actions within the dataset distribution. This approach enables the predicted actions from DT3 to be iteratively refined with Gaussian noise through the denoising chain within the diffusion model, facilitating trajectory stitching.
Figure 1 shows the performance comparison between our proposed models and the latest DT variants. Our DRDT3 significantly improves the performance of DT variants on sub-optimal datasets. Besides, DT3 without diffusion refinement also shows enhanced results over DT. The contributions of this work are summarized as follows:
@@
(iii) Experiments on extensive tasks from the D4RL benchmark [23] demonstrate the superior performance of our proposed DT3 and DRDT3 over conventional offline RL and DT-based methods.
@@
The following sections are organized as follows: we begin with preliminaries in Section II, covering offline RL, DT,the TTT layer, and diffusion models. The detailed methodology about our DRDT3 model are then presented in Section III, followed by experiments conducted for evaluating our proposed models in Section IV. Section V provides an overview of recent related work on DT, sequence modeling and diffusion models in RL. Finally, we present conclusions and discussions in Section VI.
II Preliminaries
II-A Offline Reinforcement Learning and Decision Transformer
@@
RL trains an agent by interacting with an environment that is commonly formulated as a Markov decision process (MDP): $M=\{\mathcal{S},\mathcal{A},\mathcal{R},\mathcal{P},\gamma\}$. The MDP consists of state $s\in\mathcal{S}$, action $a\in\mathcal{A}$, reward $r=\mathcal{R}(s,a)$, state transition $\mathcal{P}(s^{\prime}|s,a)$, and discount factor $\gamma\in[0,1)$ [24]. The objective of RL is to learn a policy $\pi$ that maximizes the expected cumulative discounted rewards:
@@
Different from conventional dynamic programming-based offline RL methods, DT [14] formulates the policy as a return-condition sequence model represented by GPT-like architectures [16]. Instead of generating the action based on a single given state, DT autoregressively predicts actions based on the recent context, a $K$-step sub-trajectory $\tau$ including historical tokens: states, actions, and RTGs $\hat{g}_{t}=\sum_{t^{\prime}=t}^{T}r_{t^{\prime}}$, where the RTG specifies the desired future return. The trajectory generation process can be formulated as
@@
II-B Test-Time Training Layer
Due to the quadratic complexity of self-attention in processing long-sequence inputs, some recent work leverages the RNN structure for sequence modeling, given its linear complexity. However, RNNs are often constrained by their limited expressive power in hidden states. The TTT layer [25] aims to address this challenge by updating the hidden state on both training and testing sequences with self-supervised learning. Concretely, each hidden state can be expressed as $W_{t}$, the weights of a model $f$ that produces the output $z_{t}=f(x_{t};W_{t})$. The update rule for $W_{t}$ is expressed as a step of gradient descent on a certain self-supervised loss $\ell$:
@@ II-B Test-Time Training Layer
where $\eta$ is the learning rate and $x_{t}$ is the input token. TTT layer [25] sets $\ell$ to a reconstruction loss of $x_{t}$:
@@
To make the dimensions consistent, the output rule becomes $z_{t}=f(\theta_{Q}x_{t};W_{t}).$ A sequence model with TTT layers consists of two training loops: the outer loop for updating the larger network and hyperparameters $\theta_{Q},\theta_{K}$ and $\theta_{V}$, and the inner loop for updating $W$ within each TTT layer. TTT further makes updating parallelizable through mini-batch TTT and accelerates computation using dual form technique [25].
II-C Diffusion Model
[…]
@@ II-C Diffusion Model
where $\theta$ denotes the leanable parameters of Gaussian transitions. Generating data using diffusion models involves sampling starting noises, $\mathbf{x}^{N}\sim p\left(\mathbf{x}^{N}\right)$, and implementing reverse process from timestep $i=N$ till $i=0$.
III Methodology
@@
III-A Decision TTT
The architecture of the DT3 module is depicted in the left part of Figure 2. We modify the standard DT model by replacing Transformer blocks with our proposed Attention TTT blocks for trajectory generation. The structure of the Attention TTT block is detailed in Subsection III-A1. In DT3, each input token from the context is projected into the embedding space through a linear layer, and timestep embeddings are subsequently added. Input contexts with a length smaller than $k$ are zero-padded for better generation. With the structure of DT3, actions can be generated autoregressively with known context.
III-A1 Attention TTT Block
Given the competitive performance and linear complexity of TTT layers [25] on sequence modeling, we leverage it for trajectory generating in the form of DT. Instead of using the GPT-2 model to process sequence, we propose an Attention TTT block that exploits both self-attention and the TTT layer. As shown in the left part of Figure 3, we apply masked multi-head attention to capture correlations between input tokens, followed by an Add & Norm layer. We further employ the TTT layer to extract additional information beyond what is captured by the attention mechanism. The structure of the TTT layer is illustrated in the middle part of Figure 3. In our TTT layer, we use a linear function to represent the weights $W$, resulting in $f(x)=Wx$, where $W$ is a learnable square matrix. The TTT layer is also followed by an Add & Norm layer.
III-A2 Coarse Action Prediction

### v2

@@ Abstract
Decision Transformer (DT), a trajectory modelling method, has shown competitive performance compared to traditional offline reinforcement learning (RL) approaches on various classic control tasks. However, it struggles to learn optimal policies from suboptimal, reward-labelled trajectories. In this study, we explore the use of conditional generative modelling to facilitate trajectory stitching given its high-quality data generation ability. Additionally, recent advancements in Recurrent Neural Networks (RNNs) have shown their linear complexity and competitive sequence modelling performance over Transformers. We leverage the Test-Time Training (TTT) layer, an RNN that updates hidden states during testing, to model trajectories in the form of DT. We introduce a unified framework, called Diffusion-Refined Decision TTT (DRDT3), to achieve performance beyond DT models. Specifically, we propose the Decision TTT (DT3) module, which harnesses the sequence modelling strengths of both self-attention and the TTT layer to capture recent contextual information and make coarse action predictions. DRDT3 iteratively refines the coarse action predictions through the generative diffusion model, progressively moving closer to the optimal actions. We further integrate DT3 with the diffusion model using a unified optimization objective. With experiments on multiple tasks in the D4RL benchmark, our DT3 model without diffusion refinement demonstrates improved performance over standard DT, while DRDT3 further achieves superior results compared to state-of-the-art DT-based and offline RL methods.
1 Introduction
@@
Reinforcement learning aims to train agents through interaction with the environment and receiving reward feedback. It has been widely applied in areas such as robotics (Jia et al., 2023; Guo et al., 2023), autonomous driving (Liu et al., 2024a), energy systems (Zhang et al., 2022; Bui et al., 2025), transportation systems (Zhai et al., 2025), and games (Barros e Sá & Madeira, 2025). Offline RL, also known as batch RL, eliminates the reliance on online interactions with the environment (Wang et al., 2022) by directly learning policies from offline datasets composed of historical trajectories. Therefore, it is promising for applications where interactions are risky and costly. Directly applying dynamic programming-based online RL methods to offline RL problems typically learns a poor-quality policy due to overestimated out-of-distribution actions caused by distribution shift (Yamagata et al., 2023). Conventional offline RL generally addresses such issues by regularizing policy to stay close to behavior policy (Lyu et al., 2022) or constraining values of out-of-distribution actions (Kostrikov et al., 2021; Liu et al., 2024b).
@@
DT (Chen et al., 2021), an innovative offline RL approach that formulates policy learning as a sequence modelling problem, has shown superior performance over conventional offline RL methods on some benchmarks. By leveraging the strong sequence modelling capacity of Transformer structures (Vaswani et al., 2017; Radford et al., 2019), DT models trajectories autoregressively, bypassing the necessity for bootstrapping and learning a value function for a given state. Fed with the recent context, a historical subtrajectory composed of states, actions, and return-to-go (RTG) (cumulative rewards from the current step), DT predicts actions autoregressively. However, the abandonment of dynamic programming in DT makes it lack stitching ability (Yamagata et al., 2023). Some recent work studies this problem. For instance, Q-learning DT (Yamagata et al., 2023) relabels RTG with a value function learned from Q-learning, enhancing DT to learn better policies from sub-optimal datasets. Elastic DT (Wu et al., 2023) adjusts the context length according to the evaluated optimality of the context to achieve trajectory stitching. Our work instead addresses such a problem using generative modelling.
@@
Recently, diffusion models  (Ho et al., 2020; Cao et al., 2024; Yan et al., 2024) have gained recognition for their extraordinary ability to generate high-quality complex data, such as images and texts. Additionally, recent works have also introduced the diffusion model to offline RL (Janner et al., 2022; Ajay et al., 2023; Ho et al., 2020). Similar to DT, Diffuser (Janner et al., 2022) also works as a trajectory generator, using expressive diffusion models instead of Transformers. Unlike existing diffusion-based offline RL methods, our work leverages diffusion models as a refinement tool.
@@
We propose DRDT3, a framework that combines a novel DT-style trajectory modelling method and a conditional diffusion model. Specifically, we first introduce a DT3 module, which harnesses both self-attention and the TTT layer to achieve reduced complexity and improved sequence modelling ability. This module predicts coarse action representations based on recent context, which are then used as prior knowledge or conditions for a denoising diffusion probabilistic model (DDPM) (Ho et al., 2020). Additionally, we present a novel gated MLP noise approximator, designed to capture essential information from both noisy input and coarse action predictions, facilitating the denoising process. To integrate the DT3 module and diffusion model effectively, we employ a unified optimization objective with two key components: an action representation term, which ensures that the DT3 module generates coarse action conditions close to the optimal action distribution, and an action refinement term that constrains the diffusion model to sample actions within the dataset distribution. This approach enables the predicted actions from DT3 to be iteratively refined with Gaussian noise through the denoising chain within the diffusion model, facilitating trajectory stitching.
The contributions of this work are summarized as follows:
@@
(iii) Experiments on extensive tasks from the D4RL benchmark (Fu et al., 2020) demonstrate the superior performance of our proposed DT3 and DRDT3 over conventional offline RL and DT-based methods.
@@
The following sections are organized as follows: Section 2 provides an overview of recent related work on DT, sequence modeling, and diffusion models in RL. We then present preliminaries in Section 3, covering offline RL, DT, the TTT layer, and diffusion models. Our DRDT3 methodology is detailed in Section 4, followed by experiments evaluating the proposed models in Section 5. Finally, we present conclusions and discussions in Section 6.
2 Related Work
2.1 Decision Transformer
Offline RL learns policies from historical trajectories pre-collected using other behavior policies. Different from classic offline RL (Fujimoto & Gu, 2021; Kostrikov et al., 2021) where dynamics programing is employed for policy optimization, DT (Chen et al., 2021) formulates the policy learning as a sequence modelling problem by autoregressively generating trajectories with Generative Pre-trained Transformer (GPT)-like architectures (Radford et al., 2019). Many variants of DT have been studied and shown competitive performance on some RL benchmarks. Online DT (Zheng et al., 2022) equips DT with entropy regularizers, enabling online fine-tuning with efficient exploration. Prompting DT (Xu et al., 2022) and Hyper-DT (Xu et al., 2023) make DT easily transferable to novel tasks using a trajectory prompt and an adaptation module initialized with a hyper-network, respectively. Hierarchical DT (Correia & Alexandre, 2023) frees DT from specifying RTGs while using sub-goal selection instead. Q-learning DT (Yamagata et al., 2023) and Elastic DT (Wu et al., 2023) improve the stitching ability of DT by introducing a dynamic programming-derived value function and varying-length context, respectively. DT has also been applied to some real-world applications, such as robotics (Correia & Alexandre, 2023) and traffic signal control (Huang et al., 2023).
2.2 Sequence Modeling
Transformer-based methods leveraging the attention mechanism have demonstrated remarkable performance across various sequence modelling domains (Vaswani et al., 2017; Radford et al., 2019; Achiam et al., 2023; Peebles & Xie, 2023). However, their quadratic time complexity during inference (Gu & Dao, 2023) poses challenges for long-sequence modelling. Recently, structured state space models (SSMs) (Gu et al., 2021; Gu & Dao, 2023) have introduced a compelling architecture for long-sequence processing, distinguished by their linear computational complexity. Among SSM-based methods, MAMBA has proven effective and is widely applied in various research areas (Zhang et al., 2024; Wang et al., 2024). MAMBA (Gu & Dao, 2023) introduces a selection mechanism to adapt SSM parameters based on input, along with parallel scanning, kernel fusion, and recomputation techniques for efficient computation. Other recent sequence modelling approaches, such as RWKV (Peng et al., 2023; 2024), Gated Linear Attention (GLA) (Yang et al., 2023), and xLSTM (Beck et al., 2024), build on RNN architectures with enhanced expressiveness by utilizing matrix hidden states, unlike the vector hidden states of traditional RNNs. Additionally, TTT layers (Sun et al., 2024) further enhance hidden state expressiveness by updating them through self-supervised learning during both training and inference. TTT-Linear and TTT-MLP, whose hidden states are a linear layer and an MLP, respectively, achieve linear time complexity with performance that is comparable to, or even surpasses, Transformers and MAMBA.
2.3 Diffusion Models in Reinforcement Learning
Diffusion models (Ho et al., 2020) have been appealing for their strong capacity to generate high-dimensional image or text data. In light of the expressive representation and strong multi-modal distribution modelling ability of the diffusion model, some researchers have introduced the diffusion model to RL paradigms. Diffuser (Janner et al., 2022) first employs diffusion models as a planner for generating trajectories in model-based offline RL, which alleviates the severe compounding errors of conventional planners. AdaptDiffuser (Liang et al., 2023) improves the diffusion model-based planner by self-evolving with filtered high-quality synthetic demonstrations. In addition to employing diffusion models as planners, SynthER (Lu et al., 2023), DiffStich (Li et al., 2024), GODA Huang et al. (2024), and PRIDE Feng et al. (2025) further exploit diffusion models as data synthesizers to directly augment either offline or online training data with higher diversity and quality.
Diffusion models have also been adopted to represent policies. Decision Diffuser (Ajay et al., 2023) formulates sequential decision-making problems as conditional generative modeling and introduces constraints and skill conditions. Diffusion-QL (Wang et al., 2022) explores representing policy as a diffusion model and employing Q-value function guidance during training. It overcomes the over-conservatism of policies learned from conventional offline RL. Efficient Diffusion Policy (EDP) (Kang et al., 2023) extends diffusion models to policy gradient methods and accelerates training by approximating actions from corrupted samples in a single step. Diffusion Actor-Critic (DAC) (Fang et al., 2024) adopts an actor-critic training framework, training a critic network alongside a diffusion-based KL-constrained policy, where soft Q-guidance is used to regularize the policy to remain close to the behavior policy. Ma et al. (2025) develop two diffusion-based online RL algorithms, i.e., Diffusion Policy Mirror Descent (DPMD) and Soft Diffusion Actor-Critic (SDAC), built on two novel reweighted score matching methods. MaxEntDP (Dong et al., 2025) and DIME (Celik et al., 2025) incorporate maximum entropy RL to enhance online exploration with diffusion-represented policies. DRCORL (Guo et al., 2025) applies diffusion-based policies to constrained RL and introduces a gradient operation to balance reward and safety constraints. FDPP (Chen et al., 2025) and PRIDE (Feng et al., 2025) further explore the application of diffusion models in preference-based RL. Moreover, CleanDiffuser (Dong et al., 2024) offers a diffusion model library that facilitates the development of diffusion-based decision-making methods.
Our DRDT3 uses a unified framework that aims to enhance a DT-style trajectory modeling method by refining the predictions iteratively within the expressive DDPM. In contrast to the above methods, DRDT3 employs the diffusion model solely as an action refinement module, directly operating on the DT output without multi-stage training. It remains a sequence-modeling approach without dynamic programming, avoiding the compounding errors of dynamics models and value estimation.
3 Preliminaries
3.1 Offline Reinforcement Learning and Decision Transformer
@@
RL trains an agent by interacting with an environment that is commonly formulated as a Markov decision process (MDP): $M=\{\mathcal{S},\mathcal{A},\mathcal{R},\mathcal{P},\gamma\}$. The MDP consists of state $s\in\mathcal{S}$, action $a\in\mathcal{A}$, reward $r=\mathcal{R}(s,a)$, state transition $\mathcal{P}(s^{\prime}|s,a)$, and discount factor $\gamma\in[0,1)$ (Sutton & Barto, 2018). The objective of RL is to learn a policy $\pi$ that maximizes the expected cumulative discounted rewards:
@@
Different from conventional dynamic programming-based offline RL methods, DT (Chen et al., 2021) formulates the policy as a return-condition sequence model represented by GPT-like architectures (Radford et al., 2019). Instead of generating the action based on a single given state, DT autoregressively predicts actions based on the recent context, a $K$-step sub-trajectory $\tau$ including historical tokens: states, actions, and RTGs $\hat{g}_{t}=\sum_{t^{\prime}=t}^{T}r_{t^{\prime}}$, where the RTG specifies the desired future return. The trajectory generation process can be formulated as
@@
3.2 Test-Time Training Layer
Due to the quadratic complexity of self-attention in processing long-sequence inputs, some recent work leverages the RNN structure for sequence modeling, given its linear complexity. However, RNNs are often constrained by their limited expressive power in hidden states. The TTT layer (Sun et al., 2024) aims to address this challenge by updating the hidden state on both training and testing sequences with self-supervised learning. Concretely, each hidden state can be expressed as $W_{t}$, the weights of a model $f$ that produces the output $z_{t}=f(x_{t};W_{t})$. The update rule for $W_{t}$ is expressed as a step of gradient descent on a certain self-supervised loss $\ell$:
@@ 3.2 Test-Time Training Layer
where $\eta$ is the learning rate and $x_{t}$ is the input token. TTT layer (Sun et al., 2024) sets $\ell$ to a reconstruction loss of $x_{t}$:
@@
To make the dimensions consistent, the output rule becomes $z_{t}=f(\theta_{Q}x_{t};W_{t}).$ A sequence model with TTT layers consists of two training loops: the outer loop for updating the larger network and hyperparameters $\theta_{Q},\theta_{K}$ and $\theta_{V}$, and the inner loop for updating $W$ within each TTT layer. TTT further makes updating parallelizable through mini-batch TTT and accelerates computation using the dual form technique (Sun et al., 2024).
3.3 Diffusion Model
Diffusion models (Sohl-Dickstein et al., 2015; Ho et al., 2020) typically learns expressive representation of the data distribution $q(\mathbf{x}^{0})$ from a dataset in the form $p_{\theta}\left(\mathbf{x}^{0}\right):=\int p_{\theta}\left(\mathbf{x}^{0:N}\right)d\mathbf{x}^{1:N}$, where latent variables $\mathbf{x}^{1},...,\mathbf{x}^{N}$ from timestep $i=1$ till $i=N$ have the same dimensionality as data $\mathbf{x}^{0}\sim q(\mathbf{x}^{0})$. Diffusion models typically consist of two processes: diffusion or forward process, and reverse process. The forward process is defined as a Markov chain, where the data $\mathbf{x}^{0}\sim q(\mathbf{x}^{0})$ is gradually added with Gaussian noise following a variance schedule $\beta_{1},...,\beta_{N}$:
@@ 3.3 Diffusion Model
where $\theta$ denotes the learnable parameters of Gaussian transitions. Generating data using diffusion models involves sampling starting noises, $\mathbf{x}^{N}\sim p\left(\mathbf{x}^{N}\right)$, and implementing reverse process from timestep $i=N$ till $i=0$.
4 Methodology
@@
4.1 Decision TTT
The architecture of the DT3 module is depicted in the left part of Figure 1. We modify the standard DT model by replacing Transformer blocks with our proposed Attention TTT blocks for trajectory generation. The structure of the Attention TTT block is detailed in Subsection 4.1.1. In DT3, each input token from the context is projected into the embedding space through a linear layer, and timestep embeddings are subsequently added. Input contexts with a length smaller than $k$ are zero-padded for better generation. With the structure of DT3, actions can be generated autoregressively with known context.
4.1.1 Attention TTT Block
Given the competitive performance and linear complexity of TTT layers (Sun et al., 2024) on sequence modelling, we leverage it for trajectory generation in the form of DT. Instead of using the GPT-2 model to process a sequence, we propose an Attention TTT block that exploits both self-attention and the TTT layer. As shown in the left part of Figure 2, we apply masked multi-head attention to capture correlations between input tokens, followed by an Add & Norm layer. We further employ the TTT layer to extract additional information beyond what is captured by the attention mechanism. The structure of the TTT layer is illustrated in the middle part of Figure 2. In our TTT layer, we use a linear function to represent the weights $W$, resulting in $f(x)=Wx$, where $W$ is a learnable square matrix. The TTT layer is also followed by an Add & Norm layer.
4.1.2 Coarse Action Prediction